OpenStudy (anonymous):

-5(t-4)-(t+6)=10 the solution is?

OpenStudy (anonymous):

i have -5t-20-t+6=10 for now

OpenStudy (anonymous):

wiat copied it wrong

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-5(t-4)-(t+6)=10 Distribute -5t-20-t-6=10 -6t-26=10 add liked terms together -6t=30 add 26 to the other side t=-5

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

for the last part you divide -6 to both sides

OpenStudy (anonymous):

ok so i see you added 5t to t and got 6t

OpenStudy (anonymous):

so do i have to type this as an integer or simplified fraction?